# Occlusion-Aware Consistent Model Predictive Control for Robot Navigation in Occluded Obstacle-Dense Environments

### v1

@@ Abstract
Ensuring safety and motion consistency for robot navigation in occluded, obstacle-dense environments is a critical challenge. In this context, this study presents an occlusion-aware Consistent Model Predictive Control (CMPC) strategy. To account for the occluded obstacles, it incorporates adjustable risk regions that represent their potential future locations. Subsequently, dynamic risk boundary constraints are developed online to ensure safety. The CMPC then constructs multiple locally optimal trajectory branches (each tailored to different risk regions) to balance between exploitation and exploration. A shared consensus trunk is generated to ensure smooth transitions between branches without significant velocity fluctuations, further preserving motion consistency. To facilitate high computational efficiency and ensure coordination across local trajectories, we use the alternating direction method of multipliers (ADMM) to decompose the CMPC into manageable sub-problems for parallel solving. The proposed strategy is validated through simulation and real-world experiments on an Ackermann-steering robot platform. The results demonstrate the effectiveness of the proposed CMPC strategy through comparisons with baseline approaches in occluded, obstacle-dense environments.
@@ I Introduction
Ensuring safe navigation and consistent motion for mobile robots in occluded, obstacle-dense environments is a critical challenge [1]. One of the key underlying factors to this concern is the partial observability of such environments due to occlusions [2, 3]. In this context, robots relying on onboard perception, which typically operates on line-of-sight principles, are unable to accurately detect occluded obstacles [4, 5]. This limitation increases the risk of collisions, particularly in obstacle-dense environments. Additionally, the sudden emergence of occluded obstacles can lead to abrupt velocity changes for the robot, compromising motion consistency and stability.
@@
To ensure safe navigation in obstacle-dense environments, model predictive control (MPC) facilitates obstacle avoidance through receding-horizon optimization, which systematically accounts for system dynamics and environmental constraints to achieve proactive collision avoidance [6, 7]. To construct obstacle avoidance constraints, accurate prediction of the state of the obstacle is essential for this strategy. Hence, robots must account for potential states of occluded obstacles to guarantee safety in occluded scenarios [1]. However, precise state estimation of fully occluded obstacles remains infeasible [8]. An alternative way involves predicting the probabilistic risk regions of these obstacles, which reduces planning complexity and enhances safety margins [4]. For instance, OA-MPC [1], adopts worst-case assumptions to ensure safety. This approach theoretically covers potential states of occluded obstacles. However, the generated trajectory tends to be overly conservative [9]. Moreover, a large deviation from the reference path compromises motion consistency [10].
To enhance motion consistency, recent studies have proposed scenario-adaptive planning strategies [11, 12]. In these approaches, multiple homotopic local optimal trajectories are generated for autonomous vehicles. Although consistent parameters are used within the decision-making module to select the optimal trajectory and maintain motion consistency, each trajectory is optimized separately. Essentially, the lack of coordination can lead to velocity fluctuations, particularly in occluded, obstacle-dense settings. Alternatively, branch MPC (BMPC)[13, 14] introduces a shared common trunk across branches, ensuring smooth transitions between trajectory branches and maintaining motion consistency. Similarly, partially observable Markov decision processes (POMDP)[15] ensure environmental consistency through belief state updates [16]. However, these approaches encounter scalability challenges, as branch proliferation significantly increases the number of optimization variables, leading to intractable computation. To address this challenge, recent works have employed the alternating direction method of multipliers (ADMM), which decomposes the original problem into low-dimensional sub-problems to enable real-time performance without compromising safety [17, 18, 19]. Nonetheless, despite the strong computational performance of ADMM, a more definite strategy is still lacking on how to extend its usage to address the challenges posed by occluded, obstacle-dense environments, particularly in the real-world setting.
@@
  • We introduce a consistent motion planning strategy under occlusion, where multiple trajectory branches share a common trunk as a consensus segment to ensure motion consistency. This strategy facilitates adjustable risk region configurations as dynamic risk boundary constraints in trajectory generation to strike a balance between exploitation and exploration. With this strategy, the robot exhibits less conservative behavior in occluded, obstacle-dense environments with enhanced motion consistency.
@@ II Related Work
In general, previous approaches to tackling occlusion-aware problems can be divided into reachability analysis, active perception, POMDP, and contingency planning.
@@ II-A Reachability Analysis
Reachability analysis is an efficient approach in occlusion-aware planning, as demonstrated in various studies [20, 4, 21, 22]. It creates a reachable set of potential agents in the occluded regions to construct constraints for planning. In [23], the impact of environmental visibility on mobile robot navigation is considered to ensure safety. An MPC strategy is proposed to apply risk fields from reachable sets for automated vehicles[24]. However, reachability analysis typically focuses on worst-case scenarios, leading to overly conservative behaviors. Additionally, predicting obstacle states based on rules of on-road scenarios [20, 4] is not applicable to occluded, obstacle-dense environments encountered by mobile robots.
II-B Active Perception
Active perception aims to reduce safety threats posed by potential obstacles in occluded regions. It enhances observability in the occluded regions by deviating from the reference and maximizing its field of view (FoV) in the occluded region to ensure safe navigation [25, 26, 27]. In [28, 29], the robot utilizes this approach for pursuit-evasion and tracking objects under occlusion to ensure safe navigation. However, deviating from the reference path to maximize perception is both time-consuming and energy-consuming. Additionally, excessive deviation from the reference path can significantly compromise motion consistency in unstructured and occluded, obstacle-dense environments.
II-C POMDP
POMDP has been widely adopted in occlusion-aware planning for its ability to handle partially observable scenarios [15, 30]. It leverages a probability distribution in the prediction of the agents’ future states for effective planning. For instance, works [31, 32] leverage the POMDP framework to address navigation and decision-making problems in dynamic environments with partial occlusions. [16] further incorporates a belief state updating module to predict the scenarios more precisely for effective planning. While these approaches can tackle the partially observed obstacles in structured urban road environments, their high computational complexity poses challenges for real-time planning. Additionally, obtaining reliable probability distributions is difficult in unstructured environments with dense occlusion, potentially leading to planning failures.
II-D Contingency Planning
Contingency planning enhances safety and motion consistency by providing multiple contingency paths to address uncertainties in the future states of dynamic obstacles [33, 34, 13, 35]. [36] presents a learning-based method for predicting the states of partially observed agents and motion planning on road. However, existing approaches typically require substantial computational resources to determine policies based on assumptions about future states. Additionally, contingency planning typically makes rule-based assumptions of the future states in road scenarios, or one-to-one interaction between agents [10]. These future state assumptions may not be reliable in general scenarios.
@@ III Problem Statement
$s=\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\end{bmatrix}\in\mathcal{S},\ u=\begin{bmatrix}v\\ \omega\end{bmatrix}\in\mathcal{U},$ (1)
where $p_{x}$ and $p_{y}$ denote the $x$ axis (longitudinal) and $y$ axis (lateral) position in the global coordinate system, respectively; $\theta$ denotes the heading angle of the robot; $v$ denotes the velocity of the robot; and $\omega$ denotes the changing rate of the heading angle of the robot.
@@
where $\Delta t$ denotes discrete time step; and $k$ denotes the time step. $f(s(k),u(k),\Delta t)$ is defined as follows:
@@
We make the following assumptions for this problem:
@@
The velocities of visible obstacles can be accurately detected and communicated to the robot in real time.
@@
Computational efficiency: Ensure real-time trajectory generation in occluded, obstacle-dense environments, enabling quick adaptation to dynamic conditions.
@@ IV Methodology
In this section, we propose an occlusion-aware CMPC strategy for mobile robots navigating in occluded, obstacle-dense environments. To account for occluded obstacles, we first define the occluded regions and extract risk regions. Subsequently, we formulate the motion planning problem as a CMPC. Finally, we introduce an ADMM-based optimization strategy to efficiently solve the formulated optimization problem.
@@ IV-A Risk Regions Modeling
To address the uncertainty of the states of occluded obstacles, we introduce an approach for dynamically modeling risk regions. These regions represent potential locations of occluded obstacles. The risk regions are continuously updated based on the latest relative positions of the robot and the obstacles.
@@ IV-A2 Definition of Risk Regions
Based on the defined occluded regions, we further delineate several risk regions that affect the robot’s motion. Typically, obstacles at the edge of the occluded regions pose the greatest safety threat to the robot. Therefore, these risk regions are located along the two tangent lines that form the occluded region described in Section IV-A1. Specifically, we model two circular risk regions on each tangent line, as shown in Fig. 2. For the $i$-th circular risk region, the center $(c_{x}^{(i)},c_{y}^{(i)})$ is defined as follows:
@@
By defining these risk regions, we dynamically account for the potential reachable positions of occluded obstacles. The conservatism in trajectory planning can be adjusted by making different assumptions about the maximum velocity, $v_{\text{obs,max}}$, of the occluded obstacles. In this study, we adjust $v_{\text{obs,max}}$ to generate a sequence of $N_{z}$ risk region configurations.
@@ IV-B Occlusion-Aware CMPC
This section introduces an occlusion-aware CMPC strategy with multiple branches emanating from a common trunk, referred to as a consensus segment. Each branch is configured with a distinct risk region configuration within the entire set of $N_{z}$ risk region configurations, enabling the exploration of diverse scenarios while maintaining a consensus initial trajectory segment.
@@
A set of guidance points is generated using a Visual-PRM approach [38] for each horizon. It produces optimal guidance paths and corresponding endpoints for the receding horizon based on a global task path, taking into account all visible obstacles. Specifically, guidance points provide only the endpoints of each horizon, so the corresponding cost $C_{\text{guide}}(\cdot)$ is only calculated at the time step $k=N-1$.
@@
$G_{\text{obs}}(s)=r_{\text{obs}}^{2}-\|(p_{x},p_{y})-(x_{\text{obs}},y_{{\text{obs}}})\|_{2}^{2},$
@@
$G_{\text{risk}}(s)=(r_{\text{risk}}^{(i)})^{2}-\|(p_{x},p_{y})-(c_{x}^{(i)},c_{y}^{(i)})\|_{2}^{2},$
@@ IV-C ADMM-based Optimization
To facilitate computational efficiency, we decompose the motion planning problem into several low-dimensional sub-problems. These sub-problems are solved in parallel using ADMM. While ADMM is effective in handling separable problems, the enforcement of complex constraints typically requires more sophisticated mechanisms. To address this, the augmented Lagrangian method (ALM) is employed. It integrates primal and dual variables to enforce constraints more strictly. For the inequality constraints $G_{\text{obs}}(\cdot)$ and $G_{\text{risk}}(\cdot)$, we address them by incorporating squared penalty terms [39]. For each sub-problem with the $z$-th risk region configuration, the augmented Lagrangian function is defined as follows:
@@
where $G_{\text{obs},z}(\cdot)$ and $G_{\text{risk},z}(\cdot)$ represent inequality constraints for obstacle avoidance (11b) and risk boundary (11c), respectively. $H_{\text{kin},z}(\cdot)$ represents the equality constraint related to robot kinematics (11a). $\lambda_{\text{obs},z}$, $\lambda_{\text{risk},z}$, $\lambda_{\text{kin},z}$, $\lambda_{\text{cons},z}$ are the corresponding dual variables; $\rho_{\text{obs}}$, $\rho_{\text{risk}}$, $\rho_{\text{kin}}$, $\rho_{\text{cons}}$ are the corresponding penalty coefficients. $\mathcal{I}_{+}(F)$ denotes an indicator function defined as follows:
@@
3:    $\{O^{(j)}\}_{j=0}^{N_{\text{obs}}}$ and $N_{z}$ configurations of risk regions
4:    $\{R^{(i)}_{z}\}_{z=0}^{{N_{z}}},\ i\in\mathcal{I}_{0}^{N_{\text{risk}}-1}$ via (6)-(8);
@@
6:       Reformulate into $N_{z}$ sub-problems, each with the
@@
20:    Apply next step of the consensus segment to the
21:    robot via (3);
@@
(a) Time instant $t=8\,\text{s}$.
(b) Time instant $t=10\,\text{s}$.
(c) Time instant $t=13\,\text{s}$.
Figure 3: Snapshots of the simulation, where the robot navigates in an occluded, obstacle-dense environment. The robot, shown as an orange rectangle with black wheels, optimizes three trajectories in parallel. Dark-gray boxes represent visible static obstacles, light-gray boxes represent visible dynamic obstacles, and semi-transparent boxes inside the occluded region represent the occluded obstacles from the FoV of the robot. The arrow indicates the current velocity vector of the robot. Curves in different colors represent trajectories with different risk region configurations.
[…]
Single hypothesis MPC without risk region YES $2.56$ $8.61$ $18.47$
Single hypothesis MPC with risk region NO $2.81$ $7.21$ $18.02$
CMPC NO $\mathbf{1.57}$ $\mathbf{3.71}$ $\mathbf{23.13}$
@@
For each sub-problem, (14a) is optimized using Newton’s method [40]. The dual variables corresponding to the inequality and equality constraints are updated in (14b)-(14). The optimizations of all sub-problems are performed in parallel.
Once the optimizations in (14a)-(14) of all sub-problems are complete, we update the consensus variable $\tilde{s}$ and ADMM dual variable $\lambda_{\text{cons},z}$ as follows:
$\tilde{s}^{\iota+1}=$ $\frac{1}{N_{c}}\sum_{z=0}^{N_{z}-1}s_{z}^{\iota+1},$ (15a)

### v3

@@ Abstract
Ensuring safety and motion consistency for robot navigation in occluded, obstacle-dense environments is a critical challenge. In this context, this study presents an occlusion-aware Consistent Model Predictive Control (CMPC) strategy. To account for the occluded obstacles, it incorporates adjustable risk regions that represent their potential future locations. Subsequently, dynamic risk boundary constraints are developed online to ensure safety. The CMPC then constructs multiple locally optimal trajectory branches (each tailored to different risk regions) to strike a balance between safety and performance. A shared consensus segment is generated to ensure smooth transitions between branches without significant velocity fluctuations, further preserving motion consistency. To facilitate high computational efficiency and ensure coordination across local trajectories, we use the alternating direction method of multipliers (ADMM) to decompose the CMPC into manageable sub-problems for parallel solving. The proposed strategy is validated through simulations and real-world experiments on an Ackermann-steering robot platform. The results demonstrate the effectiveness of the proposed CMPC strategy through comparisons with baseline approaches in occluded, obstacle-dense environments.
@@ I Introduction
Ensuring safe navigation and consistent motion for mobile robots in occluded, obstacle-dense environments is a critical challenge [1]. One of the key underlying factors to this concern is the partial observability of such environments due to occlusions [2, 3]. In this context, robots relying on onboard perception, which typically operate on line-of-sight principles, are unable to accurately detect occluded obstacles [4, 5]. This limitation increases the risk of collisions, particularly in obstacle-dense environments. Additionally, the sudden emergence of occluded obstacles can lead to abrupt velocity changes for the robot, compromising motion consistency and stability.
@@
To ensure safe navigation in obstacle-dense environments, model predictive control (MPC) facilitates obstacle avoidance through receding-horizon optimization, which systematically accounts for system dynamics and environmental constraints to achieve proactive collision avoidance [6, 7]. To construct obstacle avoidance constraints, accurate prediction of obstacle states is essential for this strategy. Hence, robots must account for potential states of occluded obstacles to guarantee safety in occluded scenarios [1]. However, precise state estimation of fully occluded obstacles remains infeasible [8]. An alternative way involves predicting the probabilistic risk regions of these obstacles, which reduces planning complexity and enhances safety margins [4]. For instance, OA-MPC [1] adopts worst-case assumptions to ensure safety. This approach theoretically covers potential states of occluded obstacles. However, the generated trajectory tends to be overly conservative [9]. Moreover, a large deviation from the reference path compromises motion consistency [10].
To enhance motion consistency, recent studies have proposed scenario-adaptive planning strategies [11, 12]. In these approaches, multiple homotopic local optimal trajectories are generated for autonomous vehicles. Although consistent parameters are used within the decision-making module to select the optimal trajectory and maintain motion consistency, each trajectory is optimized separately. Essentially, the lack of coordination can lead to velocity fluctuations, particularly in occluded, obstacle-dense settings. Alternatively, branch MPC (BMPC) [13, 14] introduces a shared common consensus segment across branches, ensuring smooth transitions between trajectory branches and maintaining motion consistency. Similarly, partially observable Markov decision processes (POMDP) [15] ensures environmental consistency through belief state updates [16]. However, these approaches encounter scalability challenges, as branch proliferation significantly increases the number of optimization variables, leading to intractable computation. To address this challenge, recent works have employed the alternating direction method of multipliers (ADMM), which decomposes the original problem into low-dimensional sub-problems to enable real-time performance without compromising safety [17, 18, 19]. Nonetheless, despite the strong computational performance of ADMM, a more definite strategy is still lacking on how to extend its usage to address the challenges posed by occluded, obstacle-dense environments, particularly in the real-world setting.
@@
  • We introduce a consistent motion planning strategy under occlusion, where multiple trajectory branches share a consensus segment to ensure motion consistency. This strategy facilitates adjustable risk region configurations as dynamic risk boundary constraints in trajectory generation to strike a balance between safety and performance. With this strategy, the robot exhibits less conservative behavior in occluded, obstacle-dense environments with enhanced motion consistency.
@@ II Related Work
In this section, we review three major groups of approaches to occlusion-aware planning: reachability analysis, POMDP, and contingency planning.
@@ II-A Reachability Analysis
Reachability analysis is an efficient approach in occlusion-aware planning, as demonstrated in extensive studies [20, 4, 21, 22]. It creates a reachable set of potential agents in the occluded regions to construct constraints for planning. In [23], the impact of environmental visibility on mobile robot navigation is considered to ensure safety. An MPC strategy is proposed to apply risk fields from reachable sets for automated vehicles [24]. However, reachability analysis typically focuses on worst-case scenarios, leading to overly conservative behaviors. Additionally, predicting obstacle states based on rules of on-road scenarios [20, 4] is not applicable to occluded, obstacle-dense environments encountered by mobile robots.
II-B POMDP
POMDP has been widely adopted in occlusion-aware planning for its ability to handle partially observable scenarios [15, 25]. It leverages a probability distribution in the prediction of the agents’ future states for effective planning. For instance, works [26, 27] leverage the POMDP framework to address navigation and decision-making problems in dynamic environments with partial occlusions. [16] further incorporates a belief state updating module to predict the scenarios more precisely for effective planning. While these approaches can tackle partially observed obstacles in structured urban road environments, their high computational complexity poses challenges for real-time planning. Additionally, obtaining reliable probability distributions is difficult in unstructured environments with dense occlusion, potentially leading to planning failures.
II-C Contingency Planning
Contingency planning enhances safety and motion consistency by providing multiple contingency paths to address uncertainties in the future states of dynamic obstacles [28, 29, 13, 30]. [31] presents a learning-based method for predicting the states of partially observed agents and motion planning on road. However, existing approaches typically require substantial computational resources to determine policies based on assumptions about future states. Additionally, contingency planning typically makes rule-based assumptions about the future states in road scenarios, or one-to-one interaction between agents [10]. These future state assumptions may not be reliable in general scenarios.
@@ III Problem Statement
$s=\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\end{bmatrix}\in\mathcal{S},\ u=\begin{bmatrix}v\\ \omega\end{bmatrix}\in\mathcal{U},\$(1)
where $p_{x}$ and $p_{y}$ denote the $x$-axis (longitudinal) and $y$-axis (lateral) position in the global coordinate system, respectively; $\theta$ denotes the heading angle of the robot; $v$ denotes the velocity of the robot; and $\omega$ denotes the changing rate of the heading angle of the robot.
@@
where $\Delta t$ denotes the discrete time step; and $k$ denotes the time step. $f(s(k),u(k),\Delta t)$ is defined as follows:
@@
We make the following assumption for this problem:
@@
The velocities of dynamic obstacles are bounded, and their maximum velocities are known.
@@
Computational efficiency: Ensure real-time parallel trajectory generation in occluded, obstacle-dense environments.
@@ IV Methodology
In this section, we propose an occlusion-aware CMPC strategy for mobile robots navigating in occluded, obstacle-dense environments.
@@ IV-A Risk Regions Modeling
To address the uncertainty of the states of occluded obstacles, we dynamically model risk regions to represent potential locations of occluded obstacles. The risk regions are continuously updated based on the latest states of the robot and obstacles.
@@ IV-A2 Definition of Risk Regions
Based on the defined occluded regions, we further delineate several risk regions that affect the robot’s motion. Typically, obstacles at the edge of the occluded regions pose the greatest safety threat to the robot. Therefore, these risk regions are defined along the two tangent lines that form the occluded region described in Section IV-A1. Specifically, to provide an effective geometric approximation of the polygonal occlusion cones, we model two circular risk regions on each tangent line, as shown in Fig. 2. For the $i$-th circular risk region, the center $(c_{x}^{(i)},c_{y}^{(i)})$ is defined as follows:
@@
In this study, the conservatism in trajectory planning can be adjusted by making different assumptions about the maximum velocity $v_{\text{obs, max}}$ to generate a sequence of $N_{z}$ risk region configurations.
@@
By utilizing the maximum possible obstacle velocity $v_{\text{obs,max}}$ in (8), our circular risk region model provides a conservative over-approximation of all potential positions that an occluded obstacle could occupy. This approach guarantees that the risk region encompasses any location an occluded obstacle may reach during the prediction horizon. For non-circular obstacles, we similarly use their minimum bounding circles to define the risk regions, ensuring the generality of our formulation.
@@ IV-B Occlusion-Aware CMPC
This section introduces an occlusion-aware CMPC strategy with multiple branches emanating from a common consensus segment. Each branch is configured with a distinct risk region configuration within the entire set of $N_{z}$ risk region configurations, enabling the exploration of diverse scenarios while maintaining a consensus initial trajectory segment.
@@
A set of guidance points is generated using a Visual-PRM approach [33] for each horizon. It produces optimal guidance paths for the receding horizon based on a global task path, considering all visible obstacles. Specifically, guidance points provide only the endpoints of each horizon, so the corresponding cost $C_{\text{guide}}(\cdot)$ is only calculated at the time step k=N1k=N-1.
@@
$G_{\text{obs}}(s)=r_{\text{obs}}^{2}-\|(p_{x},p_{y})-(x_{\text{obs}},y_{{\text{obs}}})\|_{2}^{2},\$
@@
$G_{\text{risk}}(s)=(r_{\text{risk}}^{(i)})^{2}-\|(p_{x},p_{y})-(c_{x}^{(i)},c_{y}^{(i)})\|_{2}^{2},\$
@@ IV-C ADMM-based Optimization
To facilitate high computational efficiency, we decompose the motion planning problem (10a)-(10e) into several low-dimensional sub-problems using the Jacobi-Proximal ADMM scheme. This method allows for independent parallel optimization of each sub-problem, with the consensus constraint ensuring synchronization across them.
While ADMM is effective in handling separable problems, the enforcement of complex constraints typically requires more sophisticated mechanisms. To address this, the augmented Lagrangian method (ALM) is employed. It integrates primal and dual variables to enforce constraints more strictly. For the inequality constraints $G_{\text{obs}}(\cdot)$ and $G_{\text{risk}}(\cdot)$, we address them by incorporating squared penalty terms [34]. For each sub-problem with the $z$-th risk region configuration, the augmented Lagrangian function is defined as follows:
@@
where $G_{\text{obs},z}(\cdot)$ and $G_{\text{risk},z}(\cdot)$ represent inequality constraints for obstacle avoidance (11b) and risk boundary (11c), respectively. $H_{\text{kin},z}(\cdot)$ represents the equality constraint related to robot kinematics (11a). $\lambda_{\text{obs},z}$, $\lambda_{\text{risk},z}$, $\lambda_{\text{kin},z}$, $\lambda_{\text{cons},z}$ are the corresponding dual variables; $\rho_{\text{obs}}$, $\rho_{\text{risk}}$, $\rho_{\text{kin}}$, $\rho_{\text{cons}}$ are the corresponding penalty coefficients. Specifically, $\rho_{\text{cons}}$ is the penalty parameter for the proximal term of the consensus constraint. $\mathcal{I}_{+}(F)$ denotes an indicator function defined as follows:
@@
[…]
@@
6:   Reformulate as $N_{z}$ sub-problems, each with the
@@
20:  Apply the first control input from the optimized
21:  consensus segment to the robot via (3);
@@
For each sub-problem, (14a) is optimized using Newton’s method [35]. The dual variables corresponding to the inequality and equality constraints are updated in (14b)-(14d). The optimizations of all sub-problems are performed in parallel.
Once the optimizations in (14a)-(14d) of all sub-problems are complete, we update the consensus variable $\tilde{s}$ and ADMM dual variable $\lambda_{\text{cons},z}$ as follows:
$\tilde{s}^{\iota+1}=$ $\frac{1}{N_{z}}\sum_{z=0}^{N_{z}-1}s_{z}^{\iota+1},$ (15a)
@@
(a) Time instant $t=8\,\text{s}$.
(b) Time instant $t=10\,\text{s}$.
(c) Time instant $t=13\,\text{s}$.
(d) Time instant $t=17\,\text{s}$.
Figure 3: Snapshots of the simulation, where the robot navigates in an occluded, obstacle-dense environment. The robot, shown as an orange rectangle with black wheels, optimizes three trajectories in parallel. Dark-gray boxes represent visible obstacles, and light-gray boxes represent the occluded obstacles from the FoV of the robot. The arrow indicates the current velocity vector of the robot. Curves in different colors represent trajectories with different risk region configurations.
Algorithm Collision Max Lat. Vel. Variance (m/s) Peak Lat. Acc. ($\text{m/s}^{2}$) Avg. Solving Time (ms)
Control-Tree YES $3.34$ $7.34$ $45.03$
Single hypothesis MPC without risk region YES $2.27$ $3.67$ $19.84$
Single hypothesis MPC with risk region NO $3.04$ $7.56$ $21.15$
CMPC-0 NO $2.55$ $7.01$ $38.01$
CMPC NO $\mathbf{1.88}$ $\mathbf{3.65}$ $\mathbf{40.78}$
TABLE II: Ablation Study Among Different Consensus Horizon Length
Consensus Horizon (s) Max Lat. Vel. Variance (m/s) Peak Lat. Acc. ($\text{m/s}^{2}$) Avg. Solving Time (ms)
0 $2.55$ $7.01$ $38.01$
1 $2.34$ $4.74$ $39.32$
2 $1.88$ $3.65$ $40.78$
5 $2.22$ $5.28$ $47.18$
Remark 4.
The computational complexity of the algorithms for each iteration can be specified as:
  • The state variable update exhibits $\mathcal{O}(N)$ complexity.
  • Averaging operations over the consensus segment scale as $\mathcal{O}(N_{c}\cdot m)$, where $N_{c}$ is the consensus horizon and $m$ is the state dimension.
  • The obstacle-related and risk region-related dual variable updates (14b) and  (14c) scale as $\mathcal{O}(N\cdot N_{\text{obs}})$ and $\mathcal{O}(N\cdot N_{\text{risk}})$, respectively. Dual variable updates (14d) and  (15b) scale as $\mathcal{O}(N)$ and $\mathcal{O}(N_{c})$, respectively.
Given that the planning horizon $N$ and consensus horizon $N_{c}$ are fixed, and the number of risk region $N_{\text{risk}}$ is bounded, the computational complexity per ADMM iteration scales linearly with the number of obstacles as $\mathcal{O}(N_{\text{obs}})$.
Remark 5.
In the Jacobi-Proximal ADMM-based optimization scheme, each sub-problem optimizes a single trajectory branch with its specific risk region configuration. In each iteration, the optimization step (14a) is solved in parallel for each risk region configuration $z$, yielding the optimal state sequence $s_{z}$ for each trajectory branch. This parallelization is key to achieving real-time performance. The solutions of all branches are then aggregated through the consensus update (15a)–(15b), which enforces that all branches share a common consensus segment, thus ensuring motion consistency. Additionally, the proximal term $\rho_{\text{cons}}|s_{z}-\tilde{s}|^{2}$ in the augmented Lagrangian (12) stabilizes the convergence of this parallel process, and this ensures numerical robustness.